An executive invests $26,000, some at 8% and some at 4% annual interest. If he receives an annual return of $1,600, how much is invested at each rate?
Answer by mananth(16946)

Part I 8.00% per annum ------------- Amount invested =x
Part II 4.00% per annum ------------ Amount invested = y
26000
Interest----- 1600.00

Part I 8.00% per annum ---x
Part II 4.00% per annum ---y
Total investment
x + 1 y= 26000 -------------1
Interest on both investments
8.00% x + 4.00% y= 1600
Multiply by 100
8 x + 4 y= 160000.00 --------2
Multiply (1) by -8
we get
-8 x -8 y= -208000.00
Add this to (2)
0 x -4 y= -48000
divide by -4
y = 12000
Part I 8.00% $ 14000
Part II 4.00% $ 12000

CHECK
14000 --------- 8.00% ------- 1120.00
12000 ------------- 4.00% ------- 480.00
Total -------------------- 1600.00

m.ananth@hotmail.ca